OpenStudy (anonymous):

A boat makes a 120-mile trip downstream in 3 hours but makes the return trip in 4 hours. What is the rate of the current? answer list 3 mph 4 mph 5 mph

OpenStudy (anonymous):

do u know how to do this

OpenStudy (paxpolaris):

using algebra Let ‘c’ be the rate of current, and ‘b’ be the rate of boat with no current...

OpenStudy (anonymous):

ok

OpenStudy (paxpolaris):

remember, distance = rate \(\times\) time \(\large D=r \cdot t\) .... or \[\implies \Large \frac D r = t\]

OpenStudy (anonymous):

40

OpenStudy (paxpolaris):

so, you get 2 equations: \[\Large {120 \over 3} = b+c\] \[\Large {120 \over 4} = b-c\] solve for c

OpenStudy (anonymous):

how do i do that

OpenStudy (paxpolaris):

have you used simultaneous equations before?

OpenStudy (paxpolaris):

did you understand how to get the two equations above?

OpenStudy (anonymous):

1 one is 40b+c an the 2 one is 30=b-c an no

OpenStudy (paxpolaris):

yea, so to get ‘c’, we subtract the second equation from the first

OpenStudy (anonymous):

how does this look like

OpenStudy (paxpolaris):

\[\large 40 = b+c\]\[\large 30 = b-c\]- - +\[\large \overline{10 = 0+2c}\]

OpenStudy (anonymous):

5=c

OpenStudy (paxpolaris):

\[\Huge \checkmark\]

OpenStudy (anonymous):

yes thx alot

OpenStudy (paxpolaris):

if you wanted to know the speed of the boat b: plug in 5 for c in any 1 of the 2 equations ...........you get 35